Please help with solving this...I can't seem to get the fractional part correct.
One number exceeds another by 5. The sum of their reciprocals equal to 19 divided by the product of the two numbers. Find the two numbers.

Answer by stanbon(48558)

Let one of the numbers be "x"; The other number is "x+5"
EQUATION:
1/x + 1/(x+5) = 19/x(x+5)
Multiply thru by x(x+5)
x+5 + x = 19
2x= 14
x = 7 (one of the numbers)
x+5 = 12 (the other number)
==============
Cheers,
Stan H.